# Liouville type theorem and Kinetic Formulation for $2\times 2$ Systems of Conservation Laws

### v1

@@ 4. Lagrangian Tools in Kinetic Setting
In the following of this Section we assume that $\boldsymbol{u}$ is an isentropic solution to (2.1) defined in $\Omega=\mathbb{R}^{+}\times\mathbb{R}$. Then, it satisfies the kinetic formulation of Proposition 3.5, i.e. it satisfies (3.20), (3.21). We assume that $\boldsymbol{u}$ is a non constant function; in particular, if $(\phi_{1},\phi_{2}):\mathcal{U}\to\mathcal{W}$ is the change of coordinates of the Riemann invariants, letting
@@
$w_{\max}\doteq\operatorname*{ess\,sup}_{t,x}w\qquad\text{for all $t\in\mathbb{R}$}$ (4.1)
$w_{\min}\doteq\operatorname*{ess\,inf}_{t,x}w\qquad\text{for all $t\in\mathbb{R}$}.$ (4.2)
@@ 5.1. VMO regularity outside of J
Let $\boldsymbol{u}:\mathbb{R}^{+}\times\mathbb{R}\to\mathcal{U}$ be a finite entropy solution (e.g. a vanishing viscosity solution when the system admits a uniformly convex entropy) to (2.1) and let $\boldsymbol{\mathrm{J}}$ be the set in (3.19). Assume that the eigenvalues are genuinely nonlinear. Then every point $(\bar{t},\bar{x})\in(0,+\infty)\times\mathbb{R}\setminus\boldsymbol{\mathrm{J}}$ is of vanishing mean oscillation, i.e.
@@
where we used (5.21), (5.22) in the first and second equality, and (4.1), (5.19) in the last inequality. ∎

### v2

@@ 4. Lagrangian Tools in Kinetic Setting
In the following of this Section we assume that $\boldsymbol{u}$ is an isentropic solution to (2.1) defined in $\Omega=\mathbb{R}^{+}\times\mathbb{R}$. Then, it satisfies the kinetic formulation of Theorem 1.3, i.e. it satisfies (1.7), (1.8) with $\mu_{i}=\nu_{i}=0$. We assume that $\boldsymbol{u}$ is a non constant function; in particular, if $(\phi_{1},\phi_{2}):\mathcal{U}\to\mathcal{W}$ is the change of coordinates of the Riemann invariants, letting
@@
$w_{\max}\doteq\operatorname*{ess\,sup}_{t,x}w$ (4.1)
$w_{\min}\doteq\operatorname*{ess\,inf}_{t,x}w$ (4.2)
@@
Recalling the definitions (3.3), (3.4) and (3.9), (3.11), notice that we have
$\mathrm{supp}\,\boldsymbol{\chi}^{\max}=\mathrm{hyp}\,\phi_{1}(\boldsymbol{u})\cap\Big{(}\mathbb{R}^{+}\times\mathbb{R}\times(w^{\max}-r,w^{\max})\Big{)}$ (4.5)
$\mathrm{supp}\,\boldsymbol{\chi}^{\min}=\mathrm{epi}\,\phi_{1}(\boldsymbol{u})\cap\Big{(}\mathbb{R}^{+}\times\mathbb{R}\times(w^{\min},w^{\min}+r)\Big{)}$ (4.6)
where $\mathrm{hyp}\,\phi_{1}(\boldsymbol{u})$ and $\mathrm{epi}\,\phi_{1}(\boldsymbol{u})$ denote the hypograph and epigraph, respectively, of the function $\phi_{1}(\boldsymbol{u})$:
$\mathrm{hyp}\,\phi_{1}(\boldsymbol{u})=\{(t,x,\xi)\;|\;\xi\leq\phi_{1}(\boldsymbol{u}(t,x))\},\qquad\mathrm{epi}\,\phi_{1}(\boldsymbol{u})=\{(t,x,\xi)\;|\;\xi\geq\phi_{1}(\boldsymbol{u}(t,x))\}$
@@ 5.1. VMO regularity outside of J
Let $\boldsymbol{u}:\mathbb{R}^{+}\times\mathbb{R}\to\mathcal{U}$ be a finite entropy solution to (2.1) and let $\boldsymbol{\mathrm{J}}$ be the set in (1.10). Assume that the eigenvalues are genuinely nonlinear. Then every point $(\bar{t},\bar{x})\in(0,+\infty)\times\mathbb{R}\setminus\boldsymbol{\mathrm{J}}$ is of vanishing mean oscillation, i.e.
@@
where we used (5.21), (5.22) in the first and second equality, and (5.19) in the last inequality. ∎